Find the equation of the line given the two points
(2,0) (0,-3)
Give each answer in y= mx+b form.
Hi I am having a lot of trouble solving these problems. Can you please explain how to solve the problem step by step?
thanks

Answer by stanbon(75887)

Find the equation of the line given the two points
(2,0) (0,-3)
Give each answer in y= mx+b form.
------------------
slope = (0--3)/(2-0) = 3/2
-----
intercept is (0,-3)
-----
Equation:
y = (3/2)x-3
===============
Cheers,
Stan H.
===============